Grade 11Physical Chemistry

The vapour density of a m ixture containing NO2 and N2O4 is 38.3 at 27`C.Calculate the no of moles of NO2 in 100 moles of mixture.

Sigma Samantasinghar
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
Vapour Density = avg. molar mass/2
av. molar mass = 2*38.3  = 74.6 = (n1*46 + n2*92)/100
46(n1 + 2*n2) = 7460
n1 + 2n2 = 7460/46 =  162 (approx.)
n1 + 2n2 = 162 …......................(1)
and given,
n1 + n2 = 100 …........................(2)
solving eqns. (1) and (2)
we get,
n2 = 62
n1 = 38